OpenStudy (bananas1234):

Solve (b + 12)^2 = 81.

rvc (rvc):

take square root on both sides

OpenStudy (bananas1234):

ok so the square root of (b +12) and the square root of 81?

rvc (rvc):

the square root of 81 is 9

OpenStudy (anonymous):

so b+12=9 b=9-12=-3 b=-3

rvc (rvc):

yep!